OpenStudy (anonymous):

Let f(x) = -3x - 4 and g(x) = 2x - 6. Find (f*g)(-7)

OpenStudy (jdoe0001):

hint: \(f(x) = -3x - 4 \qquad g(x) = 2x - 6 \\ \quad \\ (f\cdot g)(-7)\implies f(-7)\cdot g(-7)\qquad \begin{cases} f(-7)=?\\ g(-7)=? \end{cases}\)

OpenStudy (anonymous):

-340?

OpenStudy (anonymous):

A. 28 B. -20 C. 17 D. 56 These are the answers that were provided